# Audio-Conditioned Diffusion LLMs for ASR and Deliberation Processing

### v2

@@ Abstract
Diffusion-based large language models (DLLMs) have recently attracted growing interest as an alternative to autoregressive decoders. In this work, we present an empirical study on using the diffusion-based large language model LLaDA for automatic speech recognition (ASR). We first investigate its use as an external deliberation-based processing module for Whisper-LLaMA transcripts. By leveraging the bidirectional attention and denoising capabilities of LLaDA, we explore random masking, low-confidence masking, and semi-autoregressive strategies, showing that Whisper-LLaDA substantially reduces WER compared with the baseline. On LibriSpeech, the best cascade system achieves 2.25%/4.94% WER on test-clean/test-other, representing a 12.3% relative improvement over the Whisper-LLaMA baseline on the test-other split. In contrast, a plain-text LLaDA without acoustic features fails to improve accuracy, highlighting the importance of audio-conditioned embeddings. We further evaluate Whisper-LLaDA as a standalone decoder for ASR with diffusion-based and semi-autoregressive decoding. Most experimental configurations achieve faster inference than the Whisper-LLaMA baseline, although recognition accuracy is slightly lower. These findings offer an empirical view of diffusion-based LLMs for ASR and point to promising directions for improvements.
@@ 2 Diffusion-based Large Language Models
Diffusion-based large language models are defined by a generative process over discrete token sequences through a forward masking process and a learned reverse process. The forward process gradually replaces tokens in a clean sequence $x_{0}$ with a mask token [MASK] as a function of continuous time $t\in[0,1]$, where the sequence is fully revealed at given $t=0$; each token is independently masked with probability $t$ as $t$ increases, and all tokens are masked at $t=1$. The reverse process starts from a fully masked sequence and iteratively predicts masked tokens conditioned on the unmasked ones, recovering a coherent sample at $t\to 0$.
@@
where $x_{0}=[x_{0}^{1},\cdots,x_{0}^{i},\cdots,x_{0}^{L}]$ represents the clean sequence, $x_{t}$ is obtained by independently masking each position with probability $t\sim U(0,1]$, and $S_{t}$ is the set consists of the positions of all masked tokens. While the supervised fine-tuning (SFT) is implemented using the following loss function:
@@
where $(p_{0},r_{0})$ stands for paired prompt-response data, note that $p_{0}$ remains unmasked and $\mathcal{L}_{\operatorname{SFT}}$ is applied to the masked tokens of response $r_{0}$, whose positions are recorded in $S_{t}^{\prime}$. The formulation admits a time-free parameterization, the model can be trained to predict masked tokens given $x_{t}$ without taking $t$ as input, while still matching the diffusion semantics.
@@ 3 Methods
Fig. 2: Overview of decoding and deliberation-based processing strategies. (a) Diffusion-based decoding: generate the full response in parallel by iterative denoising. (b) Semi-autoregressive decoding: split the response into sub-blocks, apply diffusion within each, and proceed autoregressively across sub-blocks. (c) Diffusion-based deliberation: refine Whisper-LLaMA transcripts by randomly masking or masking low-confidence tokens, and then reconstructing them through diffusion. (d) Semi-autoregressive deliberation: refine transcripts in sub-blocks, combining diffusion within each sub-block and autoregression across sub-blocks.
@@ 3.2 Training
We trained both Whisper-LLaMA and Whisper-LLaDA on the LibriSpeech corpus [41], which contains 960 hours of English audiobooks. Speed perturbation was performed with factors of 0.9 and 1.1. During training, the Whisper encoder is frozen, while LLaMA/LLaDA are fine-tuned with LoRA. The resulting speech embeddings are concatenated with the textual prompt and ground-truth response embeddings. Following the diffusion formulation, the response block is randomly masked with probability $t\sim U(0,1]$, and LLaDA predicts the masked tokens conditioned on the remaining context. The model is optimized using cross-entropy loss over masked positions:
@@ 3.3.1 Deliberation-based Processing with LLaDA
Due to the denoising nature and bidirectional attention mechanism of the diffusion large language model, we explored a novel cascade framework for deliberation-based processing in ASR. We first obtain a transcript from the Whisper-LLaMA system to initialize the response block of Whisper-LLaDA. A proportion $p$ of response tokens are then selected and replaced with [MASK] based on the following strategies: (1) random selection; (2) select the lowest-confidence tokens, where confidence is yielded from a forward step of Whisper-LLaDA based on prompt and audio embedding; and (3) semi-autoregressive deliberation-based processing: dividing the transcript into several sub-blocks and remasking sequentially. After remasking, Whisper-LLaDA recovers the masked positions conditioned on the input prompt, the speech embeddings, and the unmasked tokens of the transcript. The reconstruction under the first two strategies is completed in one pass and the last entails one pass per sub-block, closely aligning with our training procedures. For completeness, we also attempted a plain-text version of LLaDA without Whisper embeddings.
@@ 4.1 Baseline
For reference, we also report the performance of Whisper-Large-v2/v3 (Sys. 3–4). These systems achieve very competitive results (e.g., 3.90% WER on test-other with Whisper-Large v3). We expect their larger proprietary training datasets (up to $\sim$5M hours) attribute to this. Therefore they are not directly comparable to our baselines.
Table 1: Word error rate (WER%) and real-time factor (RTF) for all proposed frameworks on LibriSpeech test-clean and test-other datasets.
@@
4.2 Deliberation-basd Processing
@@
Subsequently, we turned to Whisper-LLaDA for deliberation-based processing on Whisper-LLaMA transcripts. In this setting, we directly used the end-to-end Whisper-LLaDA checkpoint without any additional training on Whisper-LLaMA outputs. Specifically, we investigated three strategies: (1) random masking; (2) lowest-confidence tokens masking; and (3) semi-autoregressive paradigm. The results for strategies (1) and (2) are presented in Table 2, while the results for strategy (3) are summarized in Table 3.
Table 2: WER (%) of cascade deliberation-based processing with random/low-confidence masking on LibriSpeech dev/test sets. “Whisper-LLaMA” is the Whisper-LLaMA transcript baseline.
@@
Mask ratio $p$ Low-confidence masking
@@
The deliberation-based processing results demonstrate that Whisper-LLaDA yields consistent improvements over the Whisper-LLaMA baseline, confirming the effectiveness of diffusion-based refinement in recovering tokens that are challenging for autoregressive counterparts. Random masking achieves the best results when the mask ratio is 90%, while low-confidence masking brings only marginal improvements. In addition, semi-autoregressive deliberation-based processing with sub-block partition further improves recognition accuracy, suggesting that the bidirectional attention of LLaDA provides a useful supplement to the unidirectional attention in the Whisper-LLaMA model. Overall, these findings highlight the potential of Whisper-LLaDA as a complementary deliberation-based processing module for ASR systems.
@@ 4.3 Diffusion-based Decoding
We observe the following trends for diffusion decoding (Sys. 5) in Table 1. (a) Increasing the number of denoising steps steadily reduces WER, albeit with higher RTF. The best result on test-clean (2.82%) is achieved with 64 steps, while the lowest WER on test-other (5.75%) is obtained with 128 steps. These gains, however, diminish as the step count grows, indicating a saturation effect. (b) The RTFs for 1–64 steps remain consistently lower than those of AR baselines. Notably, the 64-step setting achieves a favorable trade-off, with 2.82%/5.79% WER on testclean/test-other, surpassing Sys. 2 on the latter. It also runs about $1.3\times$ faster than Sys. 1 and $2.4\times$ faster than Sys. 2, offering more efficient inference with a moderate accuracy loss relative to most AR frameworks.

### v3

# arXiv:2509.16622v3[eess.AS] 27 Feb 2026

@@ Abstract
Diffusion-based large language models (DLLMs) have recently attracted growing interest as an alternative to autoregressive decoders. In this work, we present an empirical study on using the diffusion-based large language model LLaDA for automatic speech recognition (ASR). We first investigate its use as an external deliberation-based processing module for Whisper-LLaMA transcripts. By leveraging the bidirectional attention and denoising capabilities of LLaDA, we explore random masking, lowest confidence masking, and semi-autoregressive strategies, showing that Whisper-LLaDA substantially reduces WER compared with the baseline. On LibriSpeech, the best cascade system achieves 2.25%/4.94% WER on test-clean/test-other, representing a 12.3% relative improvement over the Whisper-LLaMA baseline on the test-other split. In contrast, a plain-text LLaDA without acoustic features fails to improve accuracy, highlighting the importance of audio-conditioned embeddings. We further evaluate Whisper-LLaDA as a standalone decoder for ASR with diffusion-based and semi-autoregressive decoding. Most experimental configurations achieve faster inference than the Whisper-LLaMA baseline, although recognition accuracy is slightly lower. These findings offer an empirical view of diffusion-based LLMs for ASR and point to promising directions for improvements.
@@ 2 Diffusion-based Large Language Models
Diffusion-based large language models are defined by a generative process over discrete token sequences through a forward masking process and a learned reverse process. The forward process gradually replaces tokens in a clean sequence $x_{0}$ with a mask token [MASK] as a function of continuous time $t\in[0,1]$, where the sequence is fully observed at $t=0$; each token is independently masked with probability $t$ as $t$ increases, and all tokens are masked at $t=1$. The reverse process starts from a fully masked sequence and iteratively predicts masked tokens conditioned on the unmasked ones, recovering a coherent sample at $t\to 0$.
@@
where $x_{0}=[x_{0}^{1},\cdots,x_{0}^{i},\cdots,x_{0}^{L}]$ represents the clean sequence, $x_{t}$ is obtained by independently masking each position with probability $t\sim U(0,1]$, and $S_{t}$ is the set of positions of all masked tokens. Supervised fine-tuning (SFT) is implemented using the following loss function:
@@
where $(p_{0},r_{0})$ stands for paired prompt-response data, note that $p_{0}$ remains unmasked and $\mathcal{L}_{\operatorname{SFT}}$ is applied to the masked tokens of response $r_{0}$, whose positions are recorded in $S_{t}^{\prime}$. The formulation admits a time-free parameterization, thus the model can be trained to predict masked tokens given $x_{t}$ without taking $t$ as input, while still matching the diffusion semantics.
@@ 3 Methods
Fig. 2: Overview of decoding and deliberation-based processing strategies. (a) Diffusion-based decoding: generate the full response in parallel by iterative denoising. (b) Semi-autoregressive decoding: split the response into sub-blocks, apply diffusion within each, and proceed autoregressively across sub-blocks. (c) Diffusion-based deliberation: refine Whisper-LLaMA transcripts by randomly masking or masking lowest-confidence tokens, and then reconstructing them through diffusion. (d) Semi-autoregressive deliberation: refine transcripts in sub-blocks, combining diffusion within each sub-block and autoregression across sub-blocks.
@@ 3.2 Training
We trained both Whisper-LLaMA and Whisper-LLaDA on the LibriSpeech corpus [26], which contains 960 hours of English audiobooks. Speed perturbation was performed with factors of 0.9 and 1.1. During training, the Whisper encoder is frozen, while LLaMA/LLaDA are fine-tuned with LoRA. The resulting speech embeddings are concatenated with the textual prompt and ground-truth response embeddings. Following the diffusion formulation, tokens in the response block are independently masked with probability $t\sim U(0,1]$, and LLaDA predicts the masked tokens conditioned on the remaining context. The model is optimized using cross-entropy loss over masked positions:
@@ 3.3.1 Deliberation-based Processing with LLaDA
Due to the denoising nature and bidirectional attention mechanism of the diffusion-based large language model, we explored a novel cascade framework for deliberation-based processing in ASR. We first obtain a transcript from the Whisper-LLaMA system to initialize the response block of Whisper-LLaDA. We select a proportion $p$ of the response tokens and replace them with [MASK] based on the following strategies: (1) random selection; (2) select the lowest confidence tokens, where confidence is obtained from a forward pass of Whisper-LLaDA based on the prompt and audio embeddings; and (3) semi-autoregressive deliberation-based processing: dividing the transcript into several sub-blocks and remasking sequentially. After remasking, Whisper-LLaDA recovers the masked positions conditioned on the input prompt, the speech embeddings, and the unmasked tokens of the transcript. The reconstruction under the first two strategies is completed in steps matching the mask count and the last entails sub-block-length steps per sub-block, closely aligning with our training procedures. For completeness, we also attempted a plain-text version of LLaDA without Whisper embeddings.
@@ 4.1 Baseline
We also report the performance of Whisper-Large-v2/v3 (Sys. 3–4). These systems achieve very competitive results (e.g., 3.90% WER on test-other with Whisper-Large v3). We attribute this to their larger proprietary training datasets (up to $\sim$5M hours). Therefore, the results are provided only as reference values, as they are constructed in a manner that is not directly comparable to our systems.
Table 1: Word error rate (WER%) and real-time factor (RTF) for all proposed frameworks on LibriSpeech test-clean/other.
@@
4.2 Deliberation-based Processing
@@
Subsequently, we turned to Whisper-LLaDA for deliberation-based processing on Whisper-LLaMA transcripts. In this setting, we directly used the end-to-end Whisper-LLaDA checkpoint without any additional training on Whisper-LLaMA outputs. Specifically, we investigated three strategies: (1) random masking; (2) lowest confidence tokens masking; and (3) semi-autoregressive paradigm. The results for strategies (1) and (2) are presented in Table 2, while the results for strategy (3) are summarized in Table 3.
Table 2: WER (%) of cascade deliberation-based processing with random/lowest confidence masking on LibriSpeech dev/test sets. “Whisper-LLaMA” is the Whisper-LLaMA transcript baseline.
@@
Mask ratio $p$ Lowest confidence masking
@@
The deliberation-based processing results demonstrate that Whisper-LLaDA yields consistent improvements over the Whisper-LLaMA baseline, confirming the effectiveness of diffusion-based refinement in recovering tokens that are challenging for autoregressive counterparts. Random masking achieves the best results when the mask ratio is 90%, while lowest confidence masking brings only marginal improvements. In addition, semi-autoregressive deliberation-based processing with sub-block partition further improves recognition accuracy, suggesting that the bidirectional attention of LLaDA provides a useful supplement to the unidirectional attention in the Whisper-LLaMA model. Overall, these findings highlight the potential of Whisper-LLaDA as a complementary deliberation-based processing module for ASR systems.
@@ 4.3 Diffusion-based Decoding
We observe the following trends for diffusion decoding (Sys. 5) in Table 1. (a) Increasing the number of denoising steps steadily reduces WER, albeit with higher RTF. The best result on test-clean (2.82%) is achieved with 64 steps, while the lowest WER on test-other (5.75%) is obtained with 128 steps. These gains, however, diminish as the step count grows, indicating a saturation effect. (b) The RTFs for 1–64 steps remain consistently lower than those of AR baselines. Notably, the 64-step setting achieves a favorable trade-off, with 2.82%/5.79% WER on testclean/test-other, surpassing Sys. 2 on the latter. It operates $1.3\times$ faster than Sys. 1 and $2.4\times$ faster than Sys. 2, offering more efficient inference with a moderate accuracy loss relative to most AR frameworks.